OpenStudy (anonymous):

Use cylindrical coordinates to set up the integral of f(x, y, z) over the region outside the cone z^2=x^2+y^2 and inside x^2+y^2+z^2=9. Please don't solve it for me, I'm just having trouble finding the limits of integration.

OpenStudy (blockcolder):

Use the conversion equations from Cartesian to cylindrical and find out where the cone and the sphere intersect. :D

OpenStudy (anonymous):

Um... they intersect at x^2+y^2=9-x^2-y^2. So r=3/rad(2). Still confused. This is one of the first problems I've done of this type. Still confused on the limits.

OpenStudy (blockcolder):

I'd divide the region into two parts but it would be ugly. :(

OpenStudy (anonymous):

Help would still be appreciated.

OpenStudy (anonymous):

\[ \int _0^{\frac{3}{\sqrt{2}}}\int _r^{\sqrt{9-r^2}}rdzdr = 9-\frac{9}{\sqrt{2}} \]

OpenStudy (blockcolder):

Aren't you lacking a d(theta)?

OpenStudy (anonymous):

Multiply by 2 pi the above answer.

OpenStudy (blockcolder):

Oh yeah, My bad. =))

OpenStudy (anonymous):

\[ \int_0^{2 \pi} \int _0^{\frac{3}{\sqrt{2}}}\int _r^{\sqrt{9-r^2}}rdzdrd\theta = 2\pi \left(9-\frac{9}{\sqrt{2}}\right) \]

OpenStudy (anonymous):

You can practice similar problems and more on http://moltest.missouri.edu/mucgi-bin/calculus.cgi Chose CalcIII (Multiple Integrals)